OpenStudy (anonymous):

arctan y/x d^2f/d^2y

zepdrix (zepdrix):

You have it all jumbled up.. hard to read... \[\Large\rm f(x,y)=arcatan\left(\frac{y}{x}\right)\] And we're looking for \(\Large\rm \dfrac{d^2f}{dy^2}\), the second derivative of our function with respect to y? Or is it a partial derivative we're looking for?